A right triangle's hypotenuse has length 14. If one leg has length 8, what is the length of the other leg?



1
Expert's answer
2021-03-15T08:11:41-0400

"from \,\, pythagoras\,\, theorem\,\, ,\\
(hyp)^2 = (adj)^2 + (adj)^2 \\
\implies 14^2 = 8^2 + x^2 \\
196 = 64+x^2 \\
196 - 64= x^2 \\
132 = x^2 \\
\sqrt{132} = x \\
\therefore x = \sqrt{132} = 11.489125"